OpenStudy (goformit100):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2. y = 2x - 8 y = 2x - 4 y = 2x + 7 y = 2x + 5

OpenStudy (goformit100):

@Nurali

OpenStudy (nurali):

y-y1=m(x-x1) x1=3 and y1=-2 m=2 y+2=2(x-3) y=2(x-3)-2 y=2x-6-2 y=2x-8